OpenStudy (anonymous):

How many solutions does this system of equations have? x+y= -4 x -y= 0

OpenStudy (callisto):

x+y= -4 x -y= 0 Since slope = -a/b of the linear equation having general form ax +by+c =0 slope of equation 1 = -(1/1) = -1 slope of equation 2 = -(1/-1) = 1 Since their slope are not the same, they have a intercept. (for more info. the lines are also perpendicular to each other as the product of their slope =-1)